OpenStudy (anonymous):

Determine the integral f(x,y)dA where f(x,y)=e^(-x^2-y^2) and R is the region satisfying x>=0, y>=0, x^2+y^2<=4, and x^2+y^2>=1 Hint: use polar integration and describe the region R in terms of polar coordinates

OpenStudy (anonymous):

So first we setup the diffeomorphism between a rectangle and the annulus. $$g(r, \theta) = (r\cos \theta, r \sin \theta)$$ Now, tell me the rectangle whose image is the required set \(\{1 \leq x^2 + y^2 \leq 4\}\)

OpenStudy (anonymous):

So the restriction gives us information about the inner radius and outer radius of the annulus.

OpenStudy (anonymous):

Do you understand what we are trying to do right now? We want to find the new domain we will be using after applying the polar coordinate transformation.

OpenStudy (anonymous):

Well this is the annulus we are looking at.

OpenStudy (anonymous):

ok. so the rectangle [1,4] ?

OpenStudy (anonymous):

That's not a rectangle. Think carefully we are working in polar coordinates.

OpenStudy (anonymous):

It's \([1, 4] \times [0, 2\pi]\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok yea, all of the A's should say B correct? can you continue a bit more? Sorry.. I'm trying to understand this.

OpenStudy (anonymous):

Wait, you better double check that, there might be a mistake at the end.

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

Is 3\[\pi\] R in terms of polar coordinates?

OpenStudy (anonymous):

and that is polar coordinate terms?

OpenStudy (anonymous):

The integral is the same even though we applied the polar coordinate transform. Read about the change of variables theorem.

OpenStudy (anonymous):

ok thank you so much!

OpenStudy (anonymous):

One more correction, my apologies. So let \(A=(1,4)×(0,2\pi)\) and apply the change of variables with polar coordinate transform \(g(r,\theta)=(r \cos \theta,r \sin \theta)\) and we get $$ \newcommand{\e}{\mathrm{e}} \begin{eqnarray*} &&\int_{R}\e^{-(x^2 + y^2)} \\ &=& \int_A\e^{-((r \cos \theta)^2 + (r \sin \theta)^2)}|r| \\ &=& \int_A\e^{-r^2((\cos \theta)^2 + (\sin \theta)^2)}r \\ &=& \int_{(1, 4) \times (0, 2\pi)}\e^{-r^2}r \\ &=& \int_0^{2\pi}\int_1^4\e^{-r^2}r \\ &=& 2\pi\int_1^4\e^{-r^2}r \\ &=& 2\pi\int_1^4\e^{u}r \cdot \frac{-1}{2r}\\ &=& -\pi\int_1^4\e^{u} \\ &=& -\pi \cdot (-e+e^4) \\ &=& (\e-\e^4)\pi \end{eqnarray*} $$